OpenStudy (logan13):

Which is 216 written as a power of a number? A. 63 B. 36 C. 1082 D. 723

OpenStudy (welshfella):

Hint: divide 216 by 6

OpenStudy (bishoppatton):

4(3)3−2(−2)2+5 4(27)−2(4)+5=108−8+5=105

OpenStudy (bishoppatton):

that 3 and 2 next to that variable should have be a to the third power and b to the second power. unless the problem was really written like that.

OpenStudy (bishoppatton):

I swear the equation is 4a3−2b2+5

OpenStudy (bishoppatton):

and you just plug in a = 3 and b = -2 ... do the simple arithmetic.. that's it.